# Pathways of Thoughts: Multi-Directional Thinking for Long-form Personalized Question Answering

### v1

@@ Abstract.
Personalization is essential for adapting question answering (QA) systems to user-specific information needs, thereby improving both accuracy and user satisfaction. However, personalized QA remains relatively underexplored due to challenges such as inferring preferences from long, noisy, and implicit contexts, and generating responses that are simultaneously correct, contextually appropriate, and aligned with user expectations and background knowledge. To address these challenges, we propose Pathways of Thoughts (PoT), an inference-stage method that applies to any large language model (LLM) without requiring task-specific fine-tuning. The approach models the reasoning of an LLM as an iterative decision process, where the model dynamically selects among cognitive operations such as reasoning, revision, personalization, and clarification. This enables exploration of multiple reasoning trajectories, producing diverse candidate responses that capture different perspectives. PoT then aggregates and reweights these candidates according to inferred user preferences, yielding a final personalized response that benefits from the complementary strengths of diverse reasoning paths. Experiments on the LaMP-QA benchmark for personalized QA show that PoT consistently outperforms competitive baselines, achieving up to a 13.1% relative improvement. Human evaluation corroborates these results, with annotators preferring outputs from PoT in 66% of cases and reporting ties in only 15% of cases.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Personalizing large language models (LLMs) has emerged as an important topic due to its applications in content generation, writing assistance, education, and recommendation (Salemi et al., 2024b; Kumar et al., 2024; Mysore et al., 2023; Li et al., 2023; Naumov et al., 2019). By conditioning outputs on individual users’ preferences, contexts, and needs, personalized text generation improves user engagement, satisfaction, and overall effectiveness (Salemi et al., 2025). Personalization is especially valuable in information-seeking systems, including agentic web search and question answering–based information access (Salemi and Zamani, 2025a, b). For instance, the response to a question like “What is the best strategy to study for a difficult exam?” depends on various user-specific factors, including their background, prior knowledge, learning style, and habits. A tailored response can recommend strategies that align with the user’s cognitive strengths, subject matter, and available resources, ultimately improving the effectiveness of the guidance provided. Despite its importance, personalization in question answering remains underexplored, making it a timely and significant challenge for advancing interactive and user-centric information systems, with clear applications in web search and emerging systems such as AI mode in search. [^1] https://search.google/ways-to-search/ai-mode/
@@
Figure 1. Pathways of Thoughts overview: It explores multiple directions of thinking about answering the question, generates a response for each, and mix them into a response.
@@
Augmenting LLM inputs with personalized context of the user have proven effective for personalized text generation (Salemi et al., 2024b, a; Kumar et al., 2024; Li et al., 2023). This can be further enhanced by training the LLM to perform reasoning over the personalized context to extract the user’s preferences before response generation (Salemi et al., 2025). However, in most real-world cases, fine-tuning the LLM is often infeasible for two key reasons. On one hand, many LLMs are black boxes with frozen parameters, offering no control over their internal mechanisms. This is typical of popular applications like ChatGPT [^2] https://chat.openai.com/ and Gemini [^3] https://gemini.google.com/, which are only accessible through APIs or text interfaces. On the other hand, even with full access to model parameters, the fine-tuning process requires an inhibiting amount of computational resources for individual users. Consequently, a practical way to influence the model’s output is by applying advanced strategies during inference. Recognizing this, our paper focuses on this scenario and proposes an inference-stage strategy to achieve better personalization.
@@
Achieving effective personalization presents its own unique challenges. First, personalized question answering fundamentally differs from other QA tasks in its definition of correctness. A response must be correct from the user’s specific perspective, considering their preferences, context, and background. Even when two answers are factually accurate, one may be far more suitable for a user’s specific information needs. Second, user preferences are often not explicitly stated and must be inferred from personalized context, such as profiles or prior interactions. This context is frequently lengthy and noisy, making it difficult for a model to extract the relevant signals. These issues highlight that simply augmenting an LLM’s input with user context is often insufficient. Effective personalization requires the model to perform complex reasoning over noisy, implicit user data to deduce user preferences before generating a response.
@@
To address the aforementioned challenges of personalized question answering, this paper introduces Pathways of Thoughts (PoT), demonstrated in Figure 1. This approach, which can be applied to any LLM without additional training, enables the model to explore multiple thinking pathways when generating responses, each leading to a distinct response. We formalize each thinking pathway as a Markov Decision Process (MDP), where the LLM serves as both the agent and the environment. At each step, the LLM, acting as the agent, selects the most suitable action from a predefined set of cognitive operations, such as planning, reasoning, personalizing, and revising. The LLM then assumes the role of the environment, executing the chosen action and updating the state with newly acquired knowledge. This iterative process allows the model to refine its understanding of the user’s intent, extract key preferences from the user profile, and explore diverse strategies for answering the question. The process continues until the model determines that no further actions are needed, producing an initial set of responses. To generate the final personalized response, the LLM mixes outputs from different thinking pathways into a single response, aligning them with the user’s preferences extracted from their profile. Pathways of Thoughts addresses the challenges of LLMs in personalized question answering by allowing the model to iteratively think about the next step in responding to the question and how the lengthy user context can be effectively utilized. Additionally, it explores different diverse directions and paths, considering a range of possible solutions for answering the question. Finally, aggregation ensures that the output captures the strengths of all explored pathways, resulting in a more personalized and aligned response.
@@
We conduct our experiments on the LaMP-QA benchmark (Salemi and Zamani, 2025a), designed for evaluating personalized question answering in three diverse domains: (1) Art & Entertainment, (2) Lifestyle & Personal Development, and (3) Society & Culture. Our results demonstrate that PoT achieves the best performance across all categories, with statistically significant improvements, up to $13.1\%$ relative improvement on average compared to the baselines. Additionally, we investigate the impact of pathway length, the number of pathways, and various methods for diversifying pathways and aggregating responses on performance. We show that PoT generalizes to any long-context LLM without requiring additional training. Finally, human evaluation shows that PoT better satisfies users’ information seeking needs in $66\%$ of cases, compared to $19\%$ for the baseline.
@@ 2. Related Work
Personalization of LLMs
Personalization plays a central role in search, recommendation, and text generation (Fowler et al., 2015; Xue et al., 2009; Naumov et al., 2019; Salemi et al., 2024b). To enable personalization in LLMs, Salemi et al. (2024b) introduced a retrieval-augmented generation framework alongside the LaMP benchmark for short-form text, which was later extended to long-form content generation with LongLaMP (Kumar et al., 2024) and to personalized QA with LaMP-QA (Salemi and Zamani, 2025a). Personalized assistants have also been studied in several recent works (Li et al., 2023; Mysore et al., 2023; Lu et al., 2024; Zhang et al., 2024a). Proposed approaches span a variety of techniques, including training retrieval models with user feedback (Salemi et al., 2024a), optimizing LLMs with personalized supervision (Jang et al., 2023), and generating personalized prompts (Li et al., 2024). Parameter-efficient fine-tuning has emerged as another direction (Tan et al., 2024), including its integration with retrieval-augmented methods (Salemi and Zamani, 2024). Beyond model adaptation, reasoning and self-training strategies have been shown to improve personalization in text generation (Salemi et al., 2025). More recently, natural language feedback has proven effective for adapting systems to individual users in personalized QA (Salemi and Zamani, 2025b). Nevertheless, strategies for improving personalized QA at test time—without incurring the cost of additional training—remain underexplored, which is the focus of this work.
@@
Reasoning enables models to solve problems with step-by-step processing, known as chain-of-thought (CoT), enhancing performance in complex tasks like mathematical problem-solving, logical reasoning, and commonsense understanding (Wei et al., 2022; Liu et al., 2023; Yin et al., 2024). Training LLMs to perform intermediate reasoning steps before generating responses has been effective, as seen in GPT-o1 (OpenAI, 2024) and DeepSeek-R1 (DeepSeek-AI, 2025). Additionally, agentic and multi-turn or multi-agent systems that interact with external tools, code (Liu et al., 2025; Xie and Zou, 2024; Fu et al., 2024), simulated environments (Hao et al., 2023; Wang et al., 2024; Park et al., 2024; Yao et al., 2022), or with other agents (Guo et al., 2024) have been explored for complex problem-solving. Markov Decision Processes (MDP) have shown promise in reasoning for games and mathematical problems (Hao et al., 2023; Wang et al., 2024). However, these methods have not been applied to free-form text generation, particularly personalization which requires user-specific responses rather than objectively correct answers. This study adapts these approaches for personalized text generation, combining them with graph search instead of traditional tree search in MDPs.
@@ 3. Problem Formulation
We assume a question $x_{u}$ originates from a user $u$, who has a set of $n_{u}$ personalized information elements, denoted as $P_{u}=\{p_{i}\}_{i=1}^{n_{u}}$. Following previous work that utilizes long-term user history as the user profile (Salemi and Zamani, 2025a; Salemi et al., 2024b), $P_{u}$ consists of the user’s previously asked questions along with the their question narrative. The objective is to use this personalized information, along with the question $x_{u}$, to generate a personalized response for the user using an LLM $\pi$, formulated as $\hat{y}_{u}=\pi(x_{u},P_{u})$. To evaluate the personalized response generated for user $u$’s question, we assume access to a set of $n_{x_{u}}$ aspects $E_{x_{u}}=\{e_{i}\}_{i=1}^{n_{x_{u}}}$ that user $u$ expects to be addressed in the response, extracted from a question narrative $r_{x_{u}}$ written by the user $u$. Note that these aspects are used solely for evaluation and are not accessible to the model during response generation. Finally, a metric $\mu(x_{u},\hat{y}_{u},E_{x_{u}},r_{u})$ is used to score the generated response based on the extent to which these aspects are covered.
@@ 4. Pathways of Thoughts
To address the limitations of LLMs in personalized question answering, as discussed in Section 1, we introduce Pathways of Thoughts (PoT), as illustrated in Figure 1. In this approach, the LLM explores multiple thinking pathways to solve the problem, each leading to a separate response. We model the thinking process in each pathway as a Markov Decision Process (MDP), where the LLM functions as both the agent and the environment. At each step, the LLM, acting as the agent, selects an action from a set of fundamental cognitive actions, such as planning, reasoning, and revising. Then, the same LLM acts as the environment to execute the chosen action, updating the state with the newly acquired knowledge. This iterative process continues until no further actions are needed, leading to the generation of a final response. To produce a single personalized response based on all pathways, the LLM aggregates the generated responses from different thinking pathways, leveraging the strengths of each while aligning with the user’s preferences extracted from their profile. The details of this approach are provided in Algorithm 1. The following section elaborates on this method.
@@ 4.1. Thinking as a Markov Decision Process
The thinking process in the human mind involves leveraging and integrating a set of fundamental cognitive skills—such as reasoning, planning, and revision—that collectively form the foundation of structured thought (Frensch and Funke, 2002; Sternberg and Ben-Zeev, 2001). Drawing inspiration from sequential decision making in the human mind, we model the thinking process of LLMs as an MDP. Given the current thoughts and steps taken to generate a response, the model selects the next fundamental action from a predefined set to enhance its understanding of the question, refine its reasoning, or improve the response.
@@
To generate a personalized response to the question $x_{u}$ from user $u$, we assume the agent $\pi$ starts from an initial state $s_{0}$, which consists of the question $q_{u}$ and the user profile $P_{u}$. The agent iteratively selects the most probable action at each step using the prompt shown in Figure 7 in Appendix A, continuing this process until it either chooses a halting action (as determined by the LLM) or reaches the maximum number of allowed actions, denoted as $T$. Since our task involves text generation and lacks an external environment enforcing termination, the agent itself determines when to stop or continues until the predefined action limit is reached. Algorithm 1 (lines 3-17) demonstrates the algorithm. The following outlines the definitions of each MDP component in our framework:
@@
The state of the system is defined as the concatenation of the inputs, outputs, and decisions of the model into a sequence of tokens using a predefined prompt format. This representation captures the full context of the system like a conversation. The initial state of the system consists of the user’s question $x_{u}$, the personalized context $P_{u}$ about user $u$, and a set of instructions specifying the actions and the model’s expected behavior for each action, as shown in Figure 11 in Appendix A. Given this setup, the set of all possible states corresponds to the set of all textual sequences that this initial prompt is their prefix. As a result, the state space forms an uncountable discrete set.
@@
We consider the following fundamental actions that the agent can choose from in thinking process: answering, planning, checking for personalization, personalizing, reasoning, clarifying, summarizing, revising, and finalizing. [^4] This action, finalizing, also referred to as halting, terminates the process. The detailed definition and output format of each action is provided in Appendix A.1. In the initial state, the agent is restricted to selecting the planning action, as it must first plan the steps required to address the user’s question. In subsequent states, the agent can choose from all the actions; however, the model can only select the revising action if it has previously selected the answering action, as revision typically follows the generation of a response.
@@
Since our agent is used without training due to optimization costs, an explicit reward function is not required; we assume the agent makes reasonable choices. However, in general, an outcome reward function can be employed to assign rewards to actions based on the evaluation of the LLM’s final response to the question.

### v2

@@ Abstract.
Personalization is well studied in search and recommendation, but personalized question answering remains underexplored due to challenges in inferring preferences from long, noisy, implicit contexts and generating responses that are both accurate and aligned with user expectations. To address this, we propose Pathways of Thoughts (PoT), an inference-stage method that applies to any large language model (LLM) without task-specific fine-tuning. PoT models the thinking as an iterative decision process, where the model dynamically selects among cognitive operations such as reasoning, revision, personalization, and clarification. This enables exploration of multiple reasoning trajectories, producing diverse candidate responses that capture different perspectives. PoT then aggregates and reweights these candidates according to inferred user preferences, yielding a final personalized response that benefits from the complementary strengths of diverse reasoning paths. Experiments on the LaMP-QA benchmark show that PoT consistently outperforms competitive baselines, achieving up to a 10.8% relative improvement. Human evaluation further validates these improvements, with annotators preferring PoT in 66% of cases compared to the best-performing baseline and reporting ties in 15% of cases.
Personalization, Test-time compute scaling, Thinking LLMs
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792145isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Question answeringccs: Information systems Personalizationccs: Computing methodologies Markov decision processesccs: Computing methodologies Natural language generation
@@ 1. Introduction
Personalizing large language models (LLMs) has become increasingly important due to applications in content generation, writing assistance, education, and recommendation (Salemi et al., 2024b; Kumar et al., 2024; Mysore et al., 2023; Li et al., 2023; Naumov et al., 2019). By conditioning outputs on user preferences, context, and needs, personalization improves engagement, satisfaction, and overall effectiveness (Salemi et al., 2025). It is particularly valuable in information-seeking systems such as search engines and question answering (QA) (Salemi and Zamani, 2025a, b), where appropriate responses depend on user-specific factors like background knowledge and learning style. Despite its clear relevance to web search and emerging systems (e.g., AI-powered search modes), personalization in QA remains relatively underexplored.
@@
Figure 1. Pathways of Thoughts overview. PoT explores multiple directions of thinking about answering the question, generates a response for each, and mix them into a final response.
@@
Augmenting LLM inputs with the user’s personalized context have proven effective for personalized generation (Salemi et al., 2024b, a; Kumar et al., 2024; Li et al., 2023). This can be further enhanced by training the LLM to perform reasoning over the personalized context to extract the user’s preferences before response generation (Salemi et al., 2025). However, in most real-world cases, fine-tuning the LLM is often infeasible for two key reasons. On one hand, many LLMs are black boxes with frozen parameters, offering no control over their internal mechanisms. On the other hand, even with full access to parameters, the fine-tuning process requires substantial amount of personal data and computational resources for individual users. Consequently, a practical way to influence the model’s output is by applying advanced strategies during inference. Recognizing this, our paper focuses on this scenario and proposes an inference-stage strategy to achieve better personalization.
@@
Developing effective personalization models presents its own unique challenges. First, the quality of a response in personalized QA is determined by the user who asks the question, which is influenced by their preferences, context, and background. Therefore, even when two answers are factually accurate, one may be preferred by the user. Second, user preferences are often not explicitly stated and must be inferred from personalized context, such as profiles or prior interactions. This context is frequently lengthy and noisy, making it difficult for a model to extract relevant signals. These issues highlight that simply augmenting an LLM’s input with user context is often insufficient. Effective personalization requires the model to perform reasoning over noisy, implicit user data to deduce user preferences before generating a response.
@@
To address the aforementioned challenges of personalized QA, this paper introduces Pathways of Thoughts (PoT), demonstrated in Figure 1. This approach, which is generic and can be applied to any LLM without additional training, enables the model to explore multiple thinking pathways when generating responses, each leading to a distinct response. We formalize each thinking pathway as a Markov Decision Process (MDP), where the LLM serves as both the agent and the environment. At each step, the LLM, acting as the agent, selects the most suitable action from a predefined set of cognitive operations, such as planning, reasoning, personalizing, and revising. The LLM then assumes the role of the environment, executing the chosen action and updating the state with newly acquired knowledge. This iterative process allows the model to refine its understanding of the user’s intent, extract key preferences from the user profile, and explore diverse strategies for answering the question. The process continues until the model determines that no further actions are needed, producing an initial set of responses. To generate the final personalized response, the LLM mixes outputs from different thinking pathways into a single response, aligning them with the user’s preferences extracted from their profile. Pathways of Thoughts addresses the challenges of LLMs in personalized question answering by allowing the model to iteratively think about the next step in responding to the question and how the lengthy user context can be effectively utilized. Additionally, it explores different diverse directions and paths, considering a range of possible solutions for answering the question. Finally, generating a response by aggregating all pathways’ responses ensures that the output captures the strengths of all explored pathways, resulting in a more personalized and aligned response.
@@
We conduct our experiments on the LaMP-QA benchmark (Salemi and Zamani, 2025a)–a recent large-scale benchmark designed for evaluating personalized QA in three diverse domains: (1) Art & Entertainment, (2) Lifestyle & Personal Development, and (3) Society & Culture. Our results demonstrate that PoT outperforms competitive baselines across all categories, with up to $10.8\%$ statistically significant improvement on average compared to the best-performing baselines. Additionally, we investigate the impact of pathway length, the number of pathways, and various methods for diversifying pathways and aggregating responses on performance. We show that PoT generalizes to different long-context backbone LLMs without requiring additional training. Finally, human evaluation of comparing outputs side by side shows that PoT better satisfies users’ information needs in $66\%$ of cases, compared to $19\%$ for the best-performing baseline.
@@ 2. Related Work
LLM Personalization
Personalization plays a central role in search, recommendation, and text generation applications (Fowler et al., 2015; Xue et al., 2009; Naumov et al., 2019; Salemi et al., 2024b). To enable personalization in LLMs, Salemi et al. (2024b) introduced a retrieval-augmented generation framework alongside the LaMP benchmark for short-form text, which was later extended to long-form content generation with LongLaMP (Kumar et al., 2024) and more recently to personalized QA with LaMP-QA (Salemi and Zamani, 2025a). Personalized assistants have also been studied in several recent works (Li et al., 2023; Mysore et al., 2023; Lu et al., 2024; Zhang et al., 2024a). Proposed approaches span a variety of techniques, including training retrieval models with user feedback (Salemi et al., 2024a), optimizing LLMs with personalized supervision (Jang et al., 2023), and generating personalized prompts (Li et al., 2024). Parameter-efficient fine-tuning has emerged as another direction (Tan et al., 2024), including its integration with retrieval-augmented methods (Salemi and Zamani, 2024). Beyond model adaptation, reasoning and self-training strategies have been shown to improve personalization in text generation (Salemi et al., 2025). More recently, learning from natural language feedback has proven effective for adapting QA systems to individual users (Salemi and Zamani, 2025b). Nevertheless, strategies for improving personalized QA at test time—without incurring the cost of additional training—remain underexplored, which is the focus of this work.
@@
Reasoning enables models to solve problems with step-by-step processing, known as chain-of-thought (CoT), enhancing performance in complex tasks like mathematical problem-solving, logical reasoning, and commonsense understanding (Wei et al., 2022; Liu et al., 2023; Yin et al., 2024). Training LLMs to perform intermediate reasoning steps before generating responses has been effective, as seen in GPT-o1 (OpenAI, 2024b) and DeepSeek-R1 (DeepSeek-AI, 2025). Additionally, agentic and multi-turn or multi-agent systems that interact with external tools (Jin et al., 2025), code (Liu et al., 2025; Xie and Zou, 2024; Fu et al., 2024), simulated environments (Hao et al., 2023; Wang et al., 2024; Park et al., 2024; Yao et al., 2022), or with other agents (Guo et al., 2024) have been explored for complex problem-solving. Markov Decision Processes (MDP) have shown promise in reasoning for games and mathematical problems (Hao et al., 2023; Wang et al., 2024). However, these methods have not been applied to free-form text generation, particularly personalization which requires user-specific responses rather than objectively correct answers. This study extends these methods to personalized text generation by integrating them with graph-based search in place of traditional tree search in MDPs, enabling the model to produce responses that reflect the aggregated quality of all explored candidate solutions rather than relying on a single path.
@@ 3. Problem Formulation
We assume a question $x_{u}$ originates from a user $u$, who has a set of $n_{u}$ personalized information elements, denoted as $P_{u}=\{p_{i}\}_{i=1}^{n_{u}}$. Following previous work that utilizes long-term user history as the user profile (Salemi and Zamani, 2025a; Salemi et al., 2024b; Kasela et al., 2024), $P_{u}$ consists of the user’s previously asked questions along with the their question narrative. The objective is to use this personalized information, along with the question $x_{u}$, to generate a personalized response using an LLM $\pi$, formulated as $\hat{y}_{u}=\pi(x_{u},P_{u})$. To evaluate the personalized response generated for the user $u$’s question, following Salemi and Zamani (2025a), we assume access to a set of aspects $E_{x_{u}}=\{e_{i}\}_{i=1}^{|E_{x_{u}}|}$ that user $u$ expects to be addressed in the response, extracted from a question narrative $r_{x_{u}}$ written by the user. These aspects are used solely for evaluation and are not accessible to the model during response generation. Finally, a metric $\mu(x_{u},\hat{y}_{u},E_{x_{u}},r_{u})$ is used to score the generated response based on the extent to which these aspects are covered.
@@ 4. Pathways of Thoughts
Personalization is inherently user-specific, meaning that a response generation strategy effective for one user may not perform well for another. Consequently, relying on a single solution to generate a response for a user may not always yield the most preferred outcome for each individual user. To address this limitation, as well as those discussed in Section 1, we propose Pathways of Thoughts (PoT), illustrated in Figure 1. In this framework, the LLM explores multiple thinking pathways, each corresponding to a distinct line of thought that leads to a potential response. We model each thinking pathway as an MDP, where the LLM serves as both the agent and the environment. At each step, the LLM (acting as the agent) selects an action from a set of fundamental cognitive operations such as planning, reasoning, clarification, or revising. The same LLM (now acting as the environment) then executes the chosen action, updating the current state with newly acquired information about the question and user. This iterative process continues until convergence, producing a final response for each pathway. To generate a single personalized response, the LLM aggregates the responses obtained from different pathways, integrating their complementary strengths while aligning with the user’s preferences inferred from their profile. The complete procedure is detailed in Algorithm 1, and the next section elaborates on this method.
@@
Figure 2. Initial state and actions definition prompt used in Pathways of Thoughts.
@@ 4.1. Thinking as a Markov Decision Process
The thinking process in the human brain involves using and integrating a set of fundamental cognitive skills—e.g., reasoning, planning, and revision—that collectively form the foundation of structured thought (Frensch and Funke, 2002; Sternberg and Ben-Zeev, 2001). Inspired by theories about sequential decision making in the human brain, we model the thinking process of LLMs as an MDP. Given the current thoughts and steps taken to generate a response, the model selects the next fundamental action from a predefined set to enhance its representation of the question, refine its reasoning, and improve the generated response.
@@
To generate a personalized response for the question $x_{u}$ from the user $u$, we assume the agent $\pi$ starts from an initial state $s_{0}$, which consists of the question $q_{u}$ and the user profile $P_{u}$. The agent iteratively selects the most probable action at each step using the prompt shown in Figure 7 in Appendix A, continuing this process until it either chooses a halting action (as determined by the LLM) or reaches the maximum number of allowed actions, denoted as $T$. Since our task involves text generation and lacks an external environment enforcing termination, the agent itself determines when to stop or continues until the predefined action limit is reached. See = Algorithm 1 (lines 3-17) for details. The following subsections outlines the definitions of each MDP component in our framework:
@@
The state is defined as the concatenation of the inputs, outputs, and decisions of the model into a sequence of tokens using a predefined prompt format. This representation captures the full context of the system like a conversation. The initial state of the system consists of the user’s question $x_{u}$, the user profile $P_{u}$, and a set of instructions specifying the actions and the model’s expected behavior for each action, as shown in Figure 2. Given this setup, the set of all possible states corresponds to the set of all textual sequences that this initial prompt is their prefix. As a result, the state space forms an uncountable discrete set.
@@
We consider the following fundamental actions that the agent can choose from in thinking process: answering, planning, personalization, personalizing, reasoning, clarifying, summarizing, revising, and finalizing. [^1] This action, finalizing, also referred to as halting, terminates the process. The detailed definition of each action and output format of them is provided in Appendix A.1. In the initial state, the agent is restricted to selecting the planning action, as it must first plan the steps required to address the user’s question. In subsequent states, the agent can choose from all the actions; however, the model can only select the revising action if it has previously selected the answering action, as revision typically follows the generation of a response.
@@
Since our agent is used without training, an explicit reward function is not required; we assume the agent makes reasonable choices. However, in general, an outcome reward function can be employed to assign rewards to actions based on the evaluation of the LLM’s final response to the question.